Fill in the blanks so the left side is a perfect square trinomial. That is, complete the square.
x^2-14x____=(x-____)^2

Answer by Alan3354(69443)

Same problem, different numbers.
---------------------
Fill in the blanks so the left side is a perfect square trinomial. That is, complete the square.
x^2 + 16x = 0
-------
The form is ax^2 + bx + c
To find c and complete the square:
c = (b/2a)^2 = (16/2)^2 = 64
===========
x^2 + 16x + 64 = 0 + 64
(x + 8)^2 = 64